OpenStudy (anonymous):

How would I solve -(5/4)^2 and 5^3/3 Help please

OpenStudy (skullpatrol):

Any ideas?

OpenStudy (skullpatrol):

Failure to follow the guidelines may lead to your suspension or banning from the site. If you see any violations, please report them by using a report abuse link. OpenStudy values the Learning process - not the ‘Give you an answer’ process Don’t post only answers - guide the asker to a solution.

OpenStudy (jhannybean):

\[\large (\frac{5}{4})^2 = (\frac{5^2}{4^2}) = ?\]\[\large \frac{(5^3)}{3}= \frac{5*5*5}{3}\]

OpenStudy (anonymous):

Because notation is very important. Do have an image of the question?